Consider the following recent financials for XYZ Corporation: Income Statement Balance Sheet Sales 2,961 Assets 172,178 Debt 48,642 31,777 Equity 123,536 Costs EBIT 21,184 Taxes @ 38% 8050 Total 172,178 Total 172,178 Net Income 13,134 Assets and costs are proportional to sales. Debt and equity are not. A dividend of $3,109 was paid, and the company wishes to maintain a constant payout ratio. Next years sales are projected to grow by 27%.

1. What is the pro-forma value for equity? (Round answer to 2 decimal places. Do not round intermediate calculations. Also, do not calculate the numbers given in the income statement and balance sheet, such as the Taxes and Net income. Take them as given.).

2. What is the external financing needed using the pro-forma approach? (Round answer to 2 decimal places. Do not round intermediate calculations. Also, do not calculate the numbers given in the income statement and balance sheet, such as the Taxes and Net income. Take them as given.).

3. What is the internal growth rate? (Report answer in percentage terms and round to 2 decimal places. Do not round intermediate calculations).

4. What is the sustainable growth rate? (Report answer in percentage terms and round to 2 decimal places. Do not round intermediate calculations).
Answer #1

1. PRO-FORMA VALUE OF EQUITY.

​​ a.CurrentSales = $ 52961

b. Growth rate given = 27%

c.Expected Sales = 52961 x 1.27 = $ 67260.47

Given assets and costs are in proportion to sales

d.Expected Costs = 31777 x 1.27 = $ 40356.79

e.Expected EBIT = (67260.47 - 40356.79) = $ 26903.68

f.Tax @ 38% = 0.38 x 26903.68 = $ 10223.4

g.Expected Net Income = $ 16680.28

h.Expected Assets = 172178 x 1.27 = $ 218666.1

I.Current Dividend = $ 3109 and Current Net Income = $ 13134, Current Dividend Payout Ratio = (3109 / 13134) = 0.2367

Expecte Dividend = Constant Payout Ratio x Expected Net Income

= 0.2367 x 16680.28 = $ 3948.454

Expected Retained Earnings = RE = 16680.28 - 3948.454 = $ 12731.83

Therefore, Pro-Forma Equity Value = Existing Equity Value + RE = 123536 + 12731.83 = $ 136267.8

2.Amount of external finance needed.

New Asset Level = New Liabilities Level + New Equity Level + RE + External Financing needed

External Financing Needed = 218666.1 - 48642 - 123536 - 12731.83 = $ 33756.23

3.INTERNAL GROWTH RATE

Internal Growth Rate (Growth Rate achievable using only retained earnings) = Retention Ratio x (Net Income / ROA) x 100

= (1-Payout Ratio) x (Net Income / Total Assets) x100

= (1-0.2367) x (13134 / 172178) x 100 = 5.82 %

4.Sustainable growth rate.

Sustainable Growth Rate (Growth Rate achievable by means of external financing while mantaining the firm's existing capital structure) = (Retention Ratio) x (Net Income / Equity) x 100

= (1-0.2367) x (13134/123536) x 100 = 8.11 %